### v2

@@ 4.4.1 Impact of pre-training and fine-tuning
Table 7 shows the significance of pre-training and fine-tuning the VQ-MAE-AV model for audiovisual SER. Pre-training the model for the unmasking task on the VoxCeleb2 database substantially improves the emotion recognition performance, with the accuracy rising from $29.6$% to $81.5$%. Fine-tuning the encoder is also essential, as keeping it frozen leads to a $11$% drop in accuracy.
@@ 4.4.2 Impact of the encoder depth
Table 7 shows the impact of encoder depth on emotion recognition performance by varying the number of attention blocks ($L$), where a model with n attention blocks in the encoder will be denoted by VQ-MAE-AV-n. The results show that increasing the number of blocks in the encoder leads to improved performance to a certain extent. However, when the number of blocks becomes too high (as in the case of VQ-MAE-AV-20), there is no further improvement, and performance actually decreases by $1.1$% accuracy compared to VQ-MAE-AV-16.
4.4.3 Impact of the contrastive learning
The influence of contrastive learning on emotion recognition is illustrated in Table 7. When training VQ-MAE-AV exclusively with the contrastive loss, the achieved accuracy is $75.2$%, which is $9.1$% lower than when using only the generative loss. Notably, the combination of both losses, contrastive and generative, leads to improved emotion recognition, resulting in a $0.5$% accuracy boost compared to using the generative loss alone.
4.4.4 Impact of the encoder/decoder architecture
Table 7 shows the performance of the four different configurations of the VQ-MAE-AV model described in Section 3.5, including all possible combinations of the fusion strategies for the VQ-MAE-AV encoder and decoder: self-attention fusion for both the encoder and decoder (SAF-SAF); self-attention fusion for the encoder and cross-attention fusion for the decoder (SAF-CAF); cross-attention fusion for the encoder and self-attention fusion for the decoder (CAF-SAF); and cross-attention fusion for both the encoder and decoder (CAF-CAF). Among these configurations, CAF-CAF achieves the highest accuracy with a $1.5$% improvement over SAF-SAF, followed by CAF-SAF with a $1.3$% improvement over SAF-SAF, and then SAF-CAF with only a $0.3$% improvement. The cross-attention fusion encoder architecture achieves the best performance in emotion recognition, as shown by the CAF-CAF and CAF-SAF configurations. However, there exists a trade-off between performance and the number of model parameters. Notably, CAF-CAF involves slightly more than twice as many parameters as SAF-SAF.
4.4.5 Impact of the pooling strategy
Table 7 presents the impact of various pooling strategies when fine-tuning the model for audiovisual SER. The results reveal that attention-based pooling techniques, such as attention pooling and Query2Emo, outperform mean pooling. Among the two attention-based methods, Query2Emo outperforms attention pooling, with an accuracy gain of $2.8$%.
@@ 4.4.6 Impact of the audio and visual discrete token size
Fig. 5 shows the impact of the dimensions of the discrete tokens for the visual ($h$ and $w$) and audio ($d$) modalities on the audiovisual SER performance. The values of $h$ and $w$ represent the visual token size on the horizontal and vertical axes, respectively, while $d$ represents the token size on the frequency axis of the audio modality. Our study reveals that the performance of emotion recognition is impacted by both ($h,w$) and $d$ values. Therefore, it is important to select these parameters carefully. Based on our experiments, we recommend fixing them to ($h=w=4,~{}d=4$).
Figure 5: Impact of the discrete audio and visual token size on emotion recognition.
4.4.7 Impact of the audiovisual inputs
Table 7 compares our approach VQ-MAE-AV with its unimodal versions (VQ-MAE-V and VQ-MAE-A). The VQ-MAE-AV model exhibits improved performance over VQ-MAE-V and VQ-MAE-A, with gains of $5.6$% and $8.3$% in accuracy, respectively. This indicates the effectiveness of incorporating audiovisual information in the self-supervised paradigm, leading to better performance in emotion recognition.
@@ 5 Conclusion
Masked autoencoder modeling is a versatile self-supervised learning approach that can be adapted to various types of data. This paper introduced the VQ-MAE-AV model for learning representations of audiovisual speech data, which could be extended to other multimodal sequential data. VQ-MAE-AV took as input a discrete audio representation and a discrete visual representation obtained via two separate VQ-VAEs. These representations were then divided into multiple discrete tokens, with spatio-temporal tokens for the visual modality and spectro-temporal tokens for the audio modality. Pre-trained on the VoxCeleb2 dataset and fine-tuned on standard emotional audiovisual speech datasets, the experiments showed that the VQ-MAE-AV model effectively combined the audio and visual modalities for audiovisual SER, outperforming several state-of-the-art methods.
For future work, we plan to investigate other masking strategies and increase the resolution of the images to further improve emotion recognition performance. We could also explore using the representation learned by the proposed VQ-MAE-AV model for other downstream tasks in audiovisual speech processing, such as speaker recognition.
@@ Acknowledgments
This work was performed using computational resources from the “Mésocentre” computing center of Université Paris-Saclay, CentraleSupélec and École Normale Supérieure Paris-Saclay supported by CNRS and Région Île-de-France (https://mesocentre.universite-paris-saclay.fr/)

### v3

@@ 4.4.1 Impact of pre-training and fine-tuning
Table 8 shows the significance of pre-training and fine-tuning the VQ-MAE-AV model for audiovisual SER. Self-supervised pre-training of the model for the unmasking task on the VoxCeleb2 dataset substantially improves the emotion recognition performance, with the accuracy rising from $29.6$% to $81.5$%. Fine-tuning the encoder during the supervised training on RAVDESS is also essential, as keeping it frozen leads to a $11$% drop in accuracy.
@@ 4.4.2 Impact of the encoder depth
Table 8 shows the impact of varying the number $L$ of attention blocks in the VQ-MAE-AV encoder, in terms of emotion recognition performance (accuracy and F1 score), number of parameters, the number of floating-point operations (FLOPs) and the runtime (ms) required to make a forward pass in the model for 2 second-long input sequence. The results indicate that increasing the number of blocks in the encoder leads to improved emotion recognition performance up to $L=16$; the accuracy decreases by $1.1$% for $L=20$. Moreover, it can be seen that this increase in performance comes at the expense of a larger model that requires more FLOPs to compute the prediction.
4.4.3 Impact of the generative and contrastive loss functions
Table 8 shows the impact of using either the generative, the contrastive, or both loss functions during the self-supervised pre-training of the model. This experiment is conducted from the best-performing VQ-MAE-AV model (corresponding to the results in Section 4.3) instead of the baseline one. It can be seen that the best performance is achieved when using the two loss functions. When training VQ-MAE-AV exclusively with the contrastive (resp. generative) loss, the accuracy drops by $9.6$% (resp. $0.5$%).
4.4.4 Impact of the modality-fusion strategies at the encoder and decoder
Table 8 shows the the performance obtained with the different modality-fusion strategies described in Section 3.5, including all possible combinations for the VQ-MAE-AV encoder and decoder: self-attention fusion for both the encoder and decoder (SAF-SAF); self-attention fusion for the encoder and cross-attention fusion for the decoder (SAF-CAF); cross-attention fusion for the encoder and self-attention fusion for the decoder (CAF-SAF); and cross-attention fusion for both the encoder and decoder (CAF-CAF). As can be seen, CAF-CAF achieves the highest accuracy with a $1.5$% improvement over SAF-SAF, followed by CAF-SAF with a $1.3$% improvement over SAF-SAF, and then SAF-CAF with only a $0.3$% improvement. The cross-attention fusion architecture at the encoder achieves the best performance, as shown by the CAF-CAF and CAF-SAF configurations. However, there exists a trade-off between performance and the number of model parameters. Notably, CAF-CAF involves slightly more than twice as many parameters as SAF-SAF.
4.4.5 Impact of the emotion recognition model
Table 8 illustrates the impact of various emotion recognition models used on top of the VQ-MAE-AV encoder for supervised SER. As could be expected, attention pooling and Query2Emo outperform the naive mean pooling strategy. Among the two, Query2Emo performs the best, with an accuracy gain of $2.8$%.
4.4.6 Impact of the modalities
Table 8 compares the performance of the proposed multimodal model with the equivalent model trained using only the audio or visual modalities. It can be seen that exploiting both modalities greatly improves the performance, with accuracy gains of $5.6$% and $8.3$% compared to using only the visual and audio modalities, respectively.
@@ 4.4.7 Impact of the audio and visual discrete token size
Fig. 6 shows the impact of the dimension of the discrete visual and audio tokens in terms of emotion recognition accuracy. In this figure, $h$ and $w$ represent the horizontal and vertical dimensions of the token in the visual modality, while $d$ represents the frequency dimension of the tokens in the audio modality. Moreover, we only consider the case $h=w$. The results reveal that it is important to select these parameters carefully, as their value significantly impacts the performance. Based on our experiments, we recommend fixing them to ($h=w=4,~{}d=4$).
Fig. 6: Impact of the discrete audio and visual token size on emotion recognition.
4.4.8 Summary
The ablation study highlighted key factors that impact SER performance, including attention blocks, model structure, hyperparameters, and training losses. Fine-tuning on emotional datasets proved crucial, as VQ-MAE-AV initially learns general representations through masked token reconstruction, requiring adaptation for SER (see Table 8). Integrating both audio and visual modalities further enhanced performance over single-modality inputs (see Table 8). Additionally, increasing attention blocks in the encoder improves SER performance up to a saturation point (see Table 8). Other factors, such as cross-attention fusion, token size, contrastive learning, and emotion recognition models, offer smaller but notable improvements.
@@ 5 Conclusion
Masked autoencoding is a versatile self-supervised learning approach that can be adapted to various types of data. This paper introduced the VQ-MAE-AV model for learning representations of audiovisual speech data, which could be extended to other multimodal sequential data. VQ-MAE-AV took as input a discrete audio representation and a discrete visual representation obtained via two separate VQ-VAEs. These representations were then divided into multiple discrete tokens, with spatio-temporal tokens for the visual modality and spectro-temporal tokens for the audio modality. Pre-trained on the VoxCeleb2 dataset and fine-tuned on emotional audiovisual speech datasets, the experiments showed that the VQ-MAE-AV model effectively combines the audio and visual modalities for audiovisual SER, outperforming several state-of-the-art methods in both controlled and in-the-wild conditions.
During self-supervised pre-training, the VQ-MAE-AV model learns general-purpose audiovisual speech representations that require fine-tuning for SER. Unlike contrastive SSL approaches, where simple linear probing often yields satisfactory results, our model necessitates fine-tuning on emotional datasets to effectively adapt its representations for emotion recognition. This is a limitation, but it is also an opportunity, as the learned representations could be adapted for various tasks beyond SER (Baevski et al., 2022), such as audiovisual speech recognition, speaker identification, or cross-modal generative tasks, like in AnCoGen (Sadok et al., 2025).
VQ-MAE-AV employs masked modeling as an SSL pre-training paradigm, enabling efficient cross-modal integration. This approach allows the model to learn general audiovisual representations without relying on complex fusion mechanisms, making it adaptable to incorporate additional modalities such as human gesture and pose. This is a particularly interesting avenue for future work on multimodal human behavior analysis, in particular emotion recognition.
@@ Acknowledgments
This work was supported by Randstad corporate research chair. It was performed using computational resources from the “Mésocentre” computing center of Université Paris-Saclay, CentraleSupélec and École Normale Supérieure Paris-Saclay supported by CNRS and Région Île-de-France (https://mesocentre.universite-paris-saclay.fr/).